# Exploring the Landscape for Generative Sequence Models for Specialized Data Synthesis

### v1

@@ Abstract
Artificial Intelligence (AI) research often aims to develop models that can generalize reliably across complex datasets, yet this remains challenging in fields where data is scarce, intricate, or inaccessible. This paper introduces a novel approach that leverages three generative models of varying complexity to synthesize one of the most demanding structured datasets: Malicious Network Traffic. Our approach uniquely transforms numerical data into text, re-framing data generation as a language modeling task, which not only enhances data regularization but also significantly improves generalization and the quality of the synthetic data. Extensive statistical analyses demonstrate that our method surpasses state-of-the-art generative models in producing high-fidelity synthetic data. Additionally, we conduct a comprehensive study on synthetic data applications, effectiveness, and evaluation strategies, offering valuable insights into its role across various domains. Our code and pre-trained models are openly accessible at https://github.com/Moe-Zbeeb/Exploring-the-landscape-for-generative-models-for-specialized-data-generation.git, enabling further exploration and application of our methodology.
Index Terms—Data synthesis, machine learning, traffic generation, privacy preserving data, generative models.
@@ I Introduction
In recent years, Generative Adversarial Networks (GANs) have emerged as a powerful technique for producing realistic synthetic data (Goodfellow et al., 2014). GANs are widely applied in fields such as image generation, network traffic modeling, and healthcare data synthesis (Antoniou et al., 2018). These models replicate the statistical properties of real-world data, providing a valuable tool for augmenting datasets in cases where data is limited or sensitive (Ring et al., 2019a).
Despite their impressive results, GANs face challenges. Computational complexity and training instability have been widely documented, complicating replication across domains (Borji, 2019). Moreover, GANs’ primary focus on unstructured data raises questions about their suitability for structured numerical data, which is often critical in fields such as cybersecurity, finance, and healthcare (Ring et al., 2019a). This has fueled demand for alternative generative models capable of efficiently handling structured data while preserving the original data’s key statistical properties.
@@ 9 Packets Numeric 12
For our experiments, we converted raw network traffic data into CSV format using CICFlowmeter-V4.0, formerly known as ISCXFlowMeter. CICFlowmeter is a bi-directional flow generator and analyzer for Ethernet traffic, specifically designed for anomaly detection in cybersecurity datasets.
@@ VII Applications
Sankaranarayanan et al. proposed a generative adversarial network (GAN) designed to bridge the gap between embeddings in the learned feature space, which is instrumental in Visual Domain Adaptation (Sankaranarayanan et al., 2017). This methodology enables semantic segmentation across varied domains by using a generator to map features onto the image space, allowing the discriminator to operate effectively on these projections. The discriminator’s output serves as the basis for adversarial losses (Dong and Yang, 2019). Research has demonstrated that applying adversarial losses to the projected image space consistently outperforms applications to the feature space alone, yielding notably enhanced performance (Sankaranarayanan et al., 2018).

### v2

@@ Abstract
Artificial Intelligence (AI) research often aims to develop models that can generalize reliably across complex datasets, yet this remains challenging in fields where data is scarce, intricate, or inaccessible. This paper introduces a novel approach that leverages three generative models of varying complexity to synthesize one of the most demanding structured datasets: Malicious Network Traffic. Our approach uniquely transforms numerical data into text, re-framing data generation as a language modeling task, which not only enhances data regularization but also significantly improves generalization and the quality of the synthetic data. Extensive statistical analyses demonstrate that our method surpasses state-of-the-art generative models in producing high-fidelity synthetic data. Additionally, we conduct a comprehensive study on synthetic data applications, effectiveness, and evaluation strategies, offering valuable insights into its role across various domains. Our code and pre-trained models are openly accessible at Github, enabling further exploration and application of our methodology.
Index Terms: Data synthesis, machine learning, traffic generation, privacy preserving data, generative models.
@@ I Introduction
In recent years, Generative Adversarial Networks (GANs) have emerged as a powerful technique for producing realistic synthetic data (Goodfellow et al., 2014). GANs are widely applied in fields such as image generation, network traffic modeling, and healthcare data synthesis (Antoniou et al., 2018). These models replicate the statistical properties of real-world data, providing a valuable tool for augmenting datasets in cases where data is limited or sensitive (Ring et al., 2019b).
Despite their impressive results, GANs face challenges. Computational complexity and training instability have been widely documented, complicating replication across domains (Borji, 2019). Moreover, GANs’ primary focus on unstructured data raises questions about their suitability for structured numerical data, which is often critical in fields such as cybersecurity, finance, and healthcare (Ring et al., 2019b). This has fueled demand for alternative generative models capable of efficiently handling structured data while preserving the original data’s key statistical properties.
@@ 9 Packets Numeric 12
For our experiments, we converted raw network traffic data into CSV format using CICFlowmeter-V4.0, formerly known as ISCXFlowMeter. CICFlowmeter is a bi-directional flow generator and analyzer for Ethernet traffic, specifically designed for anomaly detection in cybersecurity datasets (Ring et al., 2019b).
@@ VII Applications
Sankaranarayanan et al. proposed a generative adversarial network (GAN) designed to bridge the gap between embeddings in the learned feature space, which is instrumental in Visual Domain Adaptation (Sankaranarayanan et al., 2017). This methodology enables semantic segmentation across varied domains by using a generator to map features onto the image space, allowing the discriminator to operate effectively on these projections. The discriminator’s output serves as the basis for adversarial losses (Sankaranarayanan et al., 2018). Research has demonstrated that applying adversarial losses to the projected image space consistently outperforms applications to the feature space alone, yielding notably enhanced performance (Sankaranarayanan et al., 2018).
@@ XI Conclusion
Acknowledgments
The authors would like to express their sincere gratitude to Dr. Samir Mustapha for his invaluable guidance and insightful contributions to the structuring and development of this paper. His support and feedback have been instrumental in enhancing the clarity and organization of our work.